OpenStudy (anonymous):

solve the system of equations by addition 3x + 2y = 27 6x - 6y =24

OpenStudy (anonymous):

multiply the first equation by 3 and add the two equations. 9x+6y=81 6x-6y=24 add 15x=105 x=7 y=3

OpenStudy (anonymous):

9x + 6y = 81 6x -6y = 24 3x = 57 x = 19 114-6y = 24 y = 15

OpenStudy (anonymous):

Oh. My bad